find dy/dx for y=a x az.....∞ It is 'a' raised to 'x' then 'a' raised to 'x' upto infinity

find dy/dx for


y=ax az.....∞


It is 'a' raised to 'x' then 'a' raised to 'x' upto infinity


 


 

Grade:11

1 Answers

bibhash jha
14 years ago

 

y=a^x^a^x .... upto inf..

=> y=a^xy

takinh log on both sides

lny=xylna

now differentiating wrt x , we get

1/y.dy/dx=xlna.dy/dx+ylna

=> dy/dx= y^2lna/(1-xylna)